OpenStudy (anonymous):

How do you simplify (x^2+x-12) / (x^2-9)

OpenStudy (anonymous):

factorization, i guess

OpenStudy (anonymous):

(x^2+x-12) = (x+4)(x-3) (x^2-9) = (x-3)(x+3) so (x^2+x-12) / (x^2-9) = (x+4)(x-3)/(x-3)(x+3) = (x+4)/(x+3)

OpenStudy (anonymous):

thanks